Theory and Applications of Categories, Vol. 20, No. 8, 2008, pp. 152–185.

KAN EXTENSIONS IN DOUBLE CATEGORIES (ON WEAK DOUBLE CATEGORIES, PART III)

MARCO GRANDIS AND ROBERT PARÉ

Abstract.

This paper deals with Kan extensions in a weak double category. Absolute Kan extensions are closely related to the orthogonal adjunctions introduced in a previous paper. The pointwise case is treated by introducing internal comma objects, which can be defined in an arbitrary double category.

Introduction

This is a sequel to two papers on the general theory of weak (or pseudo) double categories, ‘Limits in double categories’ [9] and ‘Adjoint for double categories’ [10], which will be referred to as Part I and II, respectively.

In Part I, it was proved that, in a weak double category 𝔸, all (small) double limits can be constructed from (small) products, equalisers and tabulators, the latter being the double limit of a vertical arrow. Part II deals with the natural notion of adjunctions G ⊣ R between weak double categories, where G is a colax double functor, while R is lax. It is called a colax/lax adjunction, and is an instance of an orthogonal adjunction in a particular double category (in the sense recalled here in 1.3), namely the (strict) double category 𝔻bl consisting of weak double categories, lax double functors (as horizontal arrows) and colax double functors (as vertical arrows), with suitable cells (see 1.4).

Here we address Kan extensions in a weak double category 𝔻, and we begin considering extensions for weak double categories.

Section 1 recalls some notions of Part II which are crucial here: orthogonal companions, orthogonal adjoints and 𝔻bl. Section 2 defines right Kan extensions in a pseudo double category, and studies the absolute ones. In particular, Thm. 2.5 shows that an orthogonal adjunction amounts to an absolute Kan extension of the identity.

Section 3 introduces internal comma objects, in a strict double category 𝔻; again, when this is 𝔻bl, we find the double comma F^{↓↓}R defined in Part II (for a colax F and a lax R having the same codomain). The general definition of comma objects implies that the solution is, at the same time, a horizontal and a vertical double limit, as made precise in Thm. 3.6; the comments of 3.3 give a motivation for choosing such a strong definition (which also requires 𝔻 to be strict).

Work supported by a research grant of Università di Genova.

Received by the editors 2007-04-18 and, in revised form, 2008-04-15.

Transmitted by R. F. C. Walters. Published on 2008-04-16.

2000 Mathematics Subject Classification: 8D05, 18A40, 18A25.

Key words and phrases: weak double category, weak double functor, Kan extension, adjoint functor, comma object.

\begin{matrix}c \\ ◯ \end{matrix} Marco Grandis and Robert Paré, 2008. Permission to copy for private use granted.

Section 4 defines pointwise Kan extensions, via the previous internal comma objects. Also here, an absolute Kan extension corresponding to an orthogonal adjunction is pointwise (Thm. 4.3).

Then, Section 5 studies deeper properties of comma cells. The main result is the Pasting Theorem 5.3, analogous to the pasting properties of homotopy pullbacks in algebraic topology. It follows that pointwise extensions are ‘stable under composition’ (Thm. 5.4). Kan extensions for weak double categories are briefly considered in the last two sections, taking 𝔻 to be 𝔻bl or some other double category of weak double categories. They will be the subject of Part IV.

As to literature, let us recall that strict double categories where introduced by C. Ehresmann [5, 6], while bicategories - the weak version of 2-categories - are due to J. Bénabou [1]. Weak (or pseudo) double categories, the natural common extension of these two structures, are much more recent and have probably first appeared in published form in our Part I. Other contributions on double categories, weak or strict, are referred to in the previous Parts. Furthermore, [7, 13] deal with homotopy theory and topology, [11, 3] with computer science, [12] with theoretical physics, and [2, 4, 14] with the general theory of double categories.

Size aspects (for double categories of double categories, for instance) can be easily settled working with suitable universes. A reference I.2, or I.2.3, or I.2.3.4 relates to Part I, namely its Section 2, or Subsection 2.3, or item (4) in the latter. Similarly for Part II.

1. Preliminaries

We review some points from Part II: orthogonal companions and adjoints in a double category 𝔻 (1.3), the double category 𝔻bl (1.4) and some double adjunctions (1.5).

1.1. Terminology. For double categories, we use the same terminology and notation as in the previous Parts. The composite of two horizontal arrows f: A → A′, g: A′ → A′′ is written gf, while for vertical arrows u: A ^{•} → B, v: B ^{•} → C we write u⊗v or v•u, or just vu (note the reversion and the dot-marked arrows). The boundary of a double cell α, consisting of two horizontal arrows and two vertical ones, as in the left diagram

[figure: A; f; •; u; A′; •; v; A; •; u; A; •; v; α; α; B; g; B′; B; B]  (1)

is displayed as α: (u (f, g) v) or also α: u → v. It is a special cell when the horizontal arrows are identities (as at the right). The horizontal and vertical compositions of cells are written as (α|β) and (\frac{α}{γ}); or also α | β and α⊗γ. The symbols 1_{A}, 1_{u} (resp. 1A^{•}, 1f^{•}) denote horizontal (resp. vertical) identities. The set of cells with boundary as in the left diagram above is written [u (f, g) v].

We generally work with pseudo double categories (I.7.1), also called weak double categories, where the horizontal structure behaves categorically, while the composition of vertical arrows is associative up to comparison cells α(u, v, w): (u⊗v)⊗w → u⊗(v⊗w); these are special isocells - horizontally invertible. But we always assume that vertical identities behave strictly, a useful simplification, easy to obtain.

Note that, by this unitarity assumption, the vertical composite of three arrows is well-defined whenever one of them is an identity: in this case, the associativity isocell α(u, v, w) is an identity (because of its coherence with the relevant unit-comparison cells, which are assumed to be so). Consequently, the vertical composite of three cells is welldefined whenever each of their two triples of vertical arrows falls in the previous situation: both associativity isocells are identities (even if identities are placed at different levels). We shall refer to this situation as a normal ternary (vertical) composition, of arrows or cells.

A weak double category 𝔸 contains a bicategory V𝔸 of vertical arrows and special cells, as well as (because of unitarity) a 2-category H𝔸 of horizontal arrows and ‘vertically special’ cells (I.1.9).

Now, a lax double functor R: 𝔸 → 𝕏 between pseudo double categories (II.2.1) preserves the horizontal structure in the strict sense, and the vertical one up to laxity comparisons, which are special cells (the identity and composition comparison)

R[A]: 1RA^{•} → R(1A^{•}): RA ^{•} → RA,  R[u, v]: Ru⊗Rv → R(u⊗v): RA ^{•} → RC, (2)

for A and u⊗v: A ^{•}→ B ^{•}→ C in 𝔸. All this has to satisfy naturality and coherence axioms. (To remember the direction of these cells, one can think of a vertical monad in 𝔸 as a lax double functor \underline{1} → 𝔸, defined on the singleton double category.)

This lax R is unitary if, for every A in 𝔸, the special cell R[A]: 1RA^{•} → R1A^{•} is an identity; then, by coherence, also the following cells are (for u: B ^{•} → A and v: A ^{•} → C)

R[u, 1A^{•}]: Ru⊗R1A^{•} → Ru,  R[1A^{•}, v]: R1A^{•}⊗Rv → Rv.  (3)

Dually, a colax double functor F: 𝔸 → 𝕏 has comparison cells in the opposite direction

[F A]: F (1A^{•}) → 1F A^{•},  F [u, v]: F (u⊗v) → F u⊗F v.  (4)

A pseudo (resp. strict) double functor is a lax one, whose comparison cells are horizontally invertible (resp. identities); or, equivalently, a colax one satisfying the same condition. A pseudo double functor can always be made unitary.

Finally, note that, as a consequence of (3), a unitary lax (or colax) double functor defined on the singleton \underline{1} is necessarily strict. More generally, this fact holds on every double category where all vertical compositions are trivial (i.e., all pairs of consecutive vertical arrows contain a vertical identity).

1.2. Some examples. Among the pseudo double categories listed in I.3, we shall mostly refer here to ℂat (formed of categories, functors and profunctors, I.3.1), 𝕊et (sets, mappings and spans, I.3.2) and ℝel (sets, mappings and relations, I.3.4). In ℂat, a profunctor F: A ^{•} → B is defined as a functor F: A^{op}×B → Set.

The double category 𝔻bl will be recalled below (1.4). We also consider Ehresmann’s double category of quintets ℚA on a 2-category A, where a double cell α: (F (R, S) G) is defined as a 2-cell α: GR → SF of A (cf. I.1.3); in particular ℚCat gives quintets for the 2-category of categories.

Since our attention is mostly devoted to double categories of ‘categories’ and ‘functors’, arrows in double categories will be henceforth denoted with capital letters, like R, S, T... for the horizontal arrows and F, G, H... for the vertical ones.

1.3. Orthogonal companions and adjoints. . Let us recall a few notions, from II.1.2-3. We are in a pseudo double category 𝔻.

First, the horizontal morphism R: A → B and the vertical morphism F: A ^{•} → B are presented as companions by assigning a pair (η, ϵ) of cells as below, called the unit and counit, satisfying the identities η | ϵ = 1R^{•}, η⊗ϵ = 1_{F}

[figure: A; •; 1; A; •; F; A; R; •; F; B; •; 1; η; ϵ; A; R; B; B; B]  (5)

Given R, this is equivalent (by unitarity, again) to saying that the pair (F, ϵ) satisfies the following universal property:

(a) for every cell ϵ′: (F′ (R, S) B) there is a unique cell λ: (F′ (A, S) F) such that ϵ′ = λ | ϵ

[figure: A; R; •; F′; B; •; 1; A; •; F′; A; R; •; F; B; •; 1; ϵ′; =; λ; ϵ; A′; S; B; A′; S; B; B]  (6)

(Dually, (F, η) is also characterised by a universal property, see II.1(b).) Therefore, if R has a vertical companion, this is determined up to a unique special isocell, and will often be written as R_{∗}. Companions compose in the obvious (covariant) way. Companionship is preserved by unitary lax or colax double functors.

We say that 𝔸 has vertical companions if every horizontal arrow has a vertical companion. For instance, in ℂat, the vertical companion to a functor R: A → B is the associated profunctor R_{∗}: A ^{•} → B, R_{∗}(a, b) = B(R(a), b); in 𝕊et or ℝel, the vertical companion of a function R: A → B is the associated span or relation A ^{•}→ B (often identified with R). Any double category of quintets ℚA has companions: for a map R, take the same arrow. (In fact, as proved in II.1.8, ℚA is the result of freely adding companions to the 2-category A, viewed as a double category in the usual way.)

Below, we will denote R and R_{∗} with the same letter and view it as a strong arrow, which can be placed horizontally or vertically in a diagram, as R: A → B or R: A ^{•} → B, respectively. In most relevant examples, such arrows can be identified without problem, or already are ‘the same thing’.

Secondly, transforming companionship by vertical (or horizontal) duality, the arrows R: A → B and G: B ^{•} → A are presented as orthogonal adjoints by a pair (ϵ, η) of cells as below

[figure: A; R; •; 1; B; •; G; B; •; G; B; •; 1; ϵ; η; A; A; A; R; B]  (7)

with ϵ | η = 1R^{•} and η ⊗ϵ = 1_{G}. R is called the horizontal adjoint and G the vertical one. Again, given R, these relations can be described by universal properties for (G, η) or (G, ϵ) (cf. II.1.3).

The vertical adjoint of R is determined up to a special isocell and can be written R^{∗}; vertical adjoints compose, contravariantly: (SR)^{∗} can be obtained as R^{∗}S^{∗}. 𝔸 is said to have vertical (orthogonal) adjoints if every horizontal arrow has a vertical adjoint.

In ℂat, the vertical adjoint to a functor R: A → B is the associated profunctor R^{∗}: B ^{•}→ A, defined by R^{∗}(b, a) = B(b, R(a)). In 𝕊et or ℝel, the vertical adjoint of a function R: A → B is the ‘reversed’ span or relation R^{♯}: B ^{•} → A. On the other hand, a double category of quintets ℚA does not have, generally, all vertical adjoints, since our data amount to an adjunction in the 2-category A, G ⊣ R, with ϵ: GR → 1, η: 1 → RG.

1.4. A double category of double categories. . Lax and colax double functors do not compose well. But they can be organised in a strict double category 𝔻bl (introduced in II.2.2) which was, in Part II, the basis for our general notion of an external double adjunction, between (weak) double categories, and will be, here, one of the basis for Kan extensions in the same external sense.

Its objects are pseudo double categories 𝔸, 𝔹, ...; its horizontal arrows are the lax double functors R, S... between them; its vertical arrows are colax double functors F, G... (II.2.1). A cell α

[figure: 𝔸; R; •; F; 𝔹; •; G; α; ℂ; S; 𝔻]  (8)

is - loosely speaking - a ‘horizontal transformation’ α: GR → SF (as stressed by the arrow we are placing in the square). But writing GR → SF is an abuse of notation: these composites are neither lax nor colax (just morphisms of double graphs, respecting the horizontal structure), and the coherence conditions of α require the individual knowledge of the four ‘functors’, with their comparison cells.

Precisely, the cell α consists of:

- the lax double functors R, S; the colax double functors F, G,

- maps αA: GR(A) → SF (A) and cells αu in 𝔻 (for all objects A and all vertical arrows u: A ^{•} → A′ in 𝔸),

\xymatrix{GRA \ar[rr]^{αA} \ar[dd]|{•}_{GRu} &  & SF A \ar[dd]|{•}^{SF u} \\  & αu \\ GRA′ \ar[rr]_{αA′} &  & SF A′}  (9)

These data must satisfy the naturality conditions (c0), (c1) and the coherence conditions (c2), (c3)

\begin{matrix}(c0) & αA′.GRf = SF f.αA, \\ (c1) & (GRa | αv) = (αu | SF a), \\ (c2) & (GR[A] | α1A^{•} | SF [A]) = (G[RA] | 1αA^{•} | S[F A]), \\ (c3) & (GR[u, v] | αw | SF [u, v]) = (G[Ru, Rv] | (αu⊗αv) | S[F u, F v]), \end{matrix}  (10)

for a horizontal arrow f: A → A′, a cell a: (u (f, g) v), an object A and a vertical composite w = u ⊗ v in 𝔸, respectively. (Actually, (c0) is a consequence of (c1); but writing it down makes things clearer.)

The horizontal and vertical compositions of cells, in 𝔻bl, are computed in the obvious way, based in both cases on the horizontal composition of cells (in weak double categories), which is strictly associative. This explains, in part, the fact that 𝔻bl is a strict double category, as proved in II.2.2.

1.5. External adjunctions. As we have already recalled, an orthogonal adjunction in 𝔻bl is a colax/lax adjunction G ⊣ R between weak double categories, where G is a colax double functor, while R is lax. The cells ϵ: GR → 1, η: 1 → RG of (5) will now be called the counit and unit of the adjunction. Various examples can be found in II.5. We only recall two of them.

In II.5.5, we considered the pseudo double categories 𝕊et(= 𝕊pSet) and ℂospSet of spans and cospans of sets, linked by an obvious colax/lax adjunction G ⊣ R

G: 𝕊pSet ⇄ ℂospSet :R,  ϵ: GR → 1,  η: 1 → RG,  (11)

where, loosely speaking, G turns a span into its pushout, while R turns a cospan into its pullback. Both G and R are unitary (or can be made so). One can replace Set with any category C with pullbacks and pushouts.

A variant shows a non-unitary example (II.5.6), based on a 2-category C with 2pullbacks, 2 -pushouts, comma and cocomma squares. There is then a second colax/lax adjunction C ⊣ K, where C is constructed with cocomma and K with comma squares

C: 𝕊pC ⇄ ℂospC :K,  ϵ: CK → 1,  η: 1 → KC.  (12)

2. Kan extensions in weak double categories

We introduce general and absolute Kan extensions in an arbitrary weak double category 𝔻.

2.1. Kan extensions in a 2-category. Let us recall that, in a 2-category D, the right Kan extension of an arrow S along an arrow R is an arrow G = Ran_{R}(S) equipped with a 2-cell ϵ: GR → S so that the pair (G, ϵ) is universal

[figure: I; R; J; G; J; G′; I; R; J; G′; ϵ; τ; =; α; I; S; A; A; I; S; A]  (13)

i.e., every similar pair (G′, α) factors through the former, by a unique 2-cell τ: G′ → G.

This extension is said to be absolute if, for every R′, S′, G′ as below, the following mapping between sets of 2-cells is bijective

[G′R′, S′G] → [G′R′R, S′S],  τ ↦ (ϵ | τ),  (14)
[figure: I; R; J; R′; G; J′; G′; I; R′R; J′; G′; ϵ; τ; =; α; I; S; A; S′; A′; I; S′S; A′]

(A more faithful rendering of the definition would have R′ the identity, but we stated it in this equivalent form with an eye on our version for double categories.)

On the other hand, assuming that our 2-category D has comma objects, G is the pointwise right Kan extension of S along R [15] if it is stable under comma squares: for every arrow H

[figure: H^{↓}R; P; Q; J′; H; ω; I; R; J; G; ϵ; I; S; A]  (15)

the comma-object H^{↓}R gives GH = Ran_{P} (SQ), via the pasted cell ϑ = ϵQ.Gω.

2.2. Kan extensions in a double category. We introduce now Kan extensions in a pseudo double category 𝔻. Given a vertical arrow F and two horizontal arrows R, S

[figure: I; R; •; F; J; •; G; ϵ; X; S; A]  (16)

we say that the vertical arrow G: J ^{•}→ A is the right Kan extension of F along the horizontal arrows R, S (or, from R to S), via the cell ϵ, if: (Ran) for every G′: J ^{•}→ A, the mapping [G (J, A) G′] → [F (R, S) G′], τ ↦ (ϵ | τ) is bijective.

The vertical arrow G is determined up to a special isocell and written Ran_{R,S}(F). The cell ϵ is called the counit of the extension, and a right Kan square.

Further, we say that G is an absolute extension if, for every R′, S′ and G′ as below, there is a bijection:

[G (R′, S′) G′] → [F (R′R, S′S) G′],  τ ↦ (ϵ | τ),  (17)
\xymatrix{I \ar[rr]^{R} \ar[dd]|{•}_{F} &  & J \ar[rr]^{R′} \ar[dd]|{•}^{G} &  & J′ \ar[dd]^{G′} &  & I \ar[rr]^{R′R} \ar[dd]|{•}_{F} &  & J′ \ar[dd]|{•}^{G′} \\  & ϵ &  & τ &  & = &  & α \\ X \ar[rr]_{S} &  & A \ar[rr]_{S′} &  & A′ &  & X \ar[rr]_{S′S} &  & A′}

The pointwise case needs comma objects in 𝔻 and will be treated in Section 4.

2.3. Dual notions. Applying transpose duality in 𝔻 (i.e., symmetry with respect to the ‘main diagonal’), we get a left Kan extension S = Lan_{F,G}(R): the horizontal arrow S is the left Kan extension of R along the vertical arrows F, G (or, from F to G), via the cell η, if:

\xymatrix{I \ar[rr]^{R} \ar[dd]|{•}_{F} &  & J \ar[dd]|{•}^{G} \\  & η \\ X \ar@{.>}[rr]_{S} &  & A}  (18)

(Lan) for every S′: X → A, the mapping [X (S, S′) A] → [F (R, S′) G], τ ↦ η ⊗τ is bijective.

On the other hand, horizontal duality in 𝔻 yields the notion of Kan liftings, which will not be studied here. In the left diagram below, the lifting (F, ϵ) gives a bijection τ ↦ (τ | ϵ)

[figure: I; R′; J; I; •; F′; I; R; •; F; A; G; τ; τ; ϵ; I; R; •; F; A; •; G; X; X; S; A; ϵ; X; S; A]  (19)

while, at the right hand, the lifting (R, ϵ) gives a bijection τ ↦ τ ⊗ϵ.

2.4. Theorem. [Pasting properties] Let ϵ′′ = (ϵ | ϵ′) be a horizontal composition of cells, in a pseudo double category 𝔻, and ϵ be an absolute right Kan square. Then:

[figure: I; R; •; F; J; R′; •; G; J′; R′′; •; G′; J′′; •; G′′; ϵ; ϵ′; τ; X; S; A; S′; A′; S′′; A′′]  (20)

(a) ϵ′ is a right Kan square if and only if ϵ′′ is;

(b) ϵ′ is an absolute right Kan square if and only if ϵ′′ is.

Proof. We prove (b); the proof of (a) is the same, taking R′′ and S′′ to be horizontal identities.

Suppose that ϵ′ is an absolute right Kan square; given a cell α on the ‘outer perimeter’ of (20), there is precisely one β: (G (R′′R′, S′′S′) G′′) such that (ϵ | β) = α, and then precisely one τ: (G′ (R′′, S′′) G′′) such that (ϵ′ | τ) = β.

Conversely, if (ϵ|ϵ′) is an absolute right Kan square, given a cell β: (G (R′′R′, S′′S′) G′′), the cell α = (ϵ | β) yields precisely one τ such that (ϵ | ϵ′ | τ) = (ϵ | β), which means - by the universal property of ϵ - precisely one τ such that (ϵ′ | τ) = β.

2.5. Theorem. [The formal vertical adjoint theorem] In a pseudo double category 𝔻, the horizontal arrow R has a vertical adjoint (1.3) if and only if the right Kan extension Ran_{R,1}(1_{A}) exists and is absolute. Then, G = Ran_{R,1}(1_{A}) is the vertical adjoint and the counit ϵ of the extension is also the ϵ-cell of the orthogonal adjunction

[figure: A; R; •; 1; X; •; G; ϵ; A; 1; A]  (21)

Proof. First, let R and G be orthogonal adjoints in 𝔻, with cells ϵ and η (as in the diagrams below) satisfying the equations (ϵ | η) = 1R^{•}, η⊗ϵ = 1_{G}.

Given α: (1^{•} (R′R, S′) G′), the equation (ϵ | τ) = α has precisely one solution as a cell τ: (G (R′, S′) G′), namely τ = (η | 1R′^{•})⊗α (see the left diagram below)

[figure: X; •; G; X; R′; •; 1^{•}; X′; •; 1^{•}; η; 1^{•}; A; R; •; 1^{•}; X; •; G; X; •; 1^{•}; A; R; •; 1^{•}; X; R′; X′; •; G′; ϵ; η; α; A; A; R; X; A; A; S′; A′]  (22)

Conversely, let G = Ran_{R,1}(1_{A}) be an absolute extension, via ϵ. Then, ϵ satisfies the universal property II.1.3(b), which says that it is a part of an orthogonal adjunction.

(Directly, one can obtain the other cell η by the equation (ϵ | η) = 1R^{•} in the right diagram, and then prove the second triangular identity η ⊗ ϵ = 1_{G} by cancelling ϵ in the equation (ϵ | η⊗ϵ) = ϵ. The latter follows from the previous equation by middle-four interchange.)

2.6. Examples of internal Kan extensions. . The previous statement gives examples of absolute Kan extensions in various weak double categories.

In fact, orthogonal adjoints in ℂat, 𝕊et and ℝel, have been recalled in 1.3. We have also recalled that an orthogonal adjunction in 𝔻bl is an external colax/lax adjunction G ⊣ R between weak double categories, as studied in Part II (1.5); this leads to external Kan extensions for weak double categories, whose examination will begin in Section 6.

2.7. Lemma. [Replacement] In a strict double category 𝔻, two cells ϵ, λ are given, as in the diagram below

\xymatrix{I′ \ar[rr]^{T} \ar[dd]|{•}_{L} &  & J′ \ar[dd]|{•}^{L′} \\  & λ \\ I \ar[rr]^{R} \ar[dd]|{•}_{F} &  & J \ar[dd]|{•}^{G} \\  & ϵ \\ X \ar[rr]_{S} &  & A}  (23)

(a) Suppose there exist two cells φ, μ as below, with (φ | μ⊗λ) = 1R^{•}. If GL′ = Ran_{T,S}(F L) via λ⊗ϵ, then G = Ran_{R,S}(F) via ϵ.

[figure: I; •; 1^{•}; I; R; •; M; J; •; M′; μ; φ; I′; T; •; L; J′; •; L′; λ; I; I; R; J]

(b) Suppose there exist three cells φ, μ, ψ as in the diagrams below, with 1_{L}⊗φ = 1_{L} and (λ⊗μ | ψ) = 1T^{•} . Then the converse holds: if G = Ran_{R,S}(F) via ϵ, then GL′ = Ran_{T,S}(F L) via λ⊗ϵ.

[figure: I; •; 1^{•}; I; •; M; I′; T; •; L; J′; •; L′; J′; •; 1^{•}; λ; φ; I′; •; L; I; R; •; M; J; •; M′; ψ; μ; I; I; I′; T; J′; J′]

Note. This statement will become clearer under stronger hypotheses (5.2), assuming that T is a ‘vertical deformation retract’ of R via λ, μ, φ, in the sense of 5.1.

Proof. (a) Take a ‘test’ vertical arrow G′: J ^{•}→ A and a cell α as in the left diagram below. Then, the cell λ⊗α gives a (unique) special cell ρ: (GL′ (1, 1) G′L′) such that (λ⊗ϵ | ρ) = λ⊗α; but then

(φ⊗1_{F} | μ⊗λ⊗ϵ | 1_{M′} ⊗ρ) = (φ⊗1_{F} | μ⊗λ⊗α).  (24)

Applying on both sides of this equality the hypothesis (φ | μ ⊗ λ) = 1R^{•}, we obtain (ϵ | 1_{M′}⊗ρ) = α. On the other hand, if (ϵ | τ) = (ϵ | τ′), then (λ⊗ϵ | 1_{L′}⊗τ) = (λ⊗ϵ | 1_{L′}⊗τ′) and 1_{L′}⊗τ = 1_{L′}⊗τ′, whence τ = τ′ (precomposing with 1_{M′} and recalling that L′M′ = 1).

[figure: I′; •; L; I′; T; •; J′; •; L′; J′; •; 1^{•}; I; •; 1^{•}; I; R; •; M; J; •; M′; 1_{L}; λ; μ; I; •; 1^{•}; I; R; •; M; J; •; M′; ψ; φ; I′; T; •; L; J′; •; L′; μ; λ; φ; I′; T; •; L; J′; •; K; J′; •; K; I; •; F; I; R; •; F; J; •; G′; 1_{F}; α; I; •; F; I; •; F; β; 1_{K}; X; X; S; A; 1_{F}; X; X; S; A; A]  (25)

(b) Now G = Ran_{R,S}(F) via ϵ, and we must prove that GL′ = Ran_{T,S}(F L), via λ⊗ϵ. Take a cell β as in the right diagram above. Then, for the cell (φ⊗1_{F} | μ⊗β) there exists a (unique) cell τ such that (ϵ | τ) = (φ⊗1_{F} | μ⊗β); therefore, the cell (1_{L′} ⊗τ | ψ ⊗1_{K}) satisfies

(λ⊗ϵ | 1_{L′}⊗τ | ψ⊗1_{K}) = (λ⊗(φ⊗1_{F} | μ⊗β) | ψ⊗1_{K}) = (1_{L}⊗φ⊗1_{F} | ((λ⊗μ | ψ)⊗β)) = β, (26)

because (λ ⊗ μ | ψ) = 1 and 1_{L} ⊗ φ = 1. Finally, if (λ ⊗ ϵ | ρ) = (λ ⊗ ϵ | ρ′), we have (μ⊗λ⊗ϵ | ρ) = (μ⊗λ⊗ϵ | ρ′) and, precomposing with φ⊗1_{F}, (ϵ | ρ) = (ϵ | ρ′), whence ρ = ρ′.

3. Internal comma objects

Comma objects in 𝔻bl have been constructed in Part II, and shown to satisfy both a horizontal and a vertical universal property. Here, we define comma objects in an abstract (strict) double category 𝔻, by a system of three universal properties based on the orthogonal companions of 1.3 and the commutative cells they produce. (The last notion makes sense also in weak double categories.)

3.1. Commutative cells. In the weak double category 𝔻 four arrows are given, as below

[figure: •; R; •; F; •; •; G; •; S; •]  (27)

In the following two cases one can define the commutative cell on this boundary, using a ‘flipping procedure’ already considered in II.1.6 in a more general form. It will be written as λ: (GR = SF); in a diagram, such commutative cells will often be denoted by a boldface Greek letter.

(a) If the vertical arrows F, G are strong (with horizontal companions denoted by the same letters) and GR = SF, then λ is defined by the pasting of the left diagram below (containing the unit of G, the counit of F and two vertical identities)

[figure: •; R; •; 1^{•}; •; •; 1^{•}; •; •; G_{∗}; •; •; F; •; R; •; F; •; R; •; •; 1^{•}; 1^{•}; η_{G}; 1_{F}; ϵ_{R}; •; R; F; •; F_{∗}; \begin{matrix}• \\ • \end{matrix}; G; S; •; 1^{•}; •; •; 1^{•}; •; •; 1^{•}; _{_{•} _{•}}; •; S; •; G; •; •; G; ϵ_{F}; 1^{•}; η_{S}; 1_{G}; •; •; S; •; •; S; •; •]  (28)

(b) If the horizontal arrows R, S are strong (with vertical companions denoted by the same letters) and SF = GR (i.e., S_{∗}F = GR_{∗}), then λ is defined by the pasting of the right diagram above.

If all the four arrows of the boundary are strong, we can proceed either way, and we get strictly the same result, provided we proceed in a coherent way, as made explicit in Lemma 3.5. Here, one should carefully distinguish between horizontal and vertical arrows, starting with a commutative square GR = SF of horizontal arrows (for instance), then replacing F and G in diagram (27) with their vertical companions F_{∗} and G_{∗}, and constructing, with the given units and counits, a special isocell realising the identification G_{∗}R_{∗} = S_{∗}F_{∗}.

Let us assume we have chosen in 𝔻 a ‘coherent choice of companions’, consistent with composition (also for units and counits); this choice need not be global, but might be restricted to a suitable subcategory of horizontal arrows. Then, with reference to this system, a commutative cell is determined by its boundary and commutative cells are closed under horizontal and vertical composition (forming a flat double subcategory of 𝔻).

3.2. Comma objects. Suppose we have, in the strict double category 𝔻, a vertical arrow F and a horizontal arrow R with the same codomain. The (internal) comma object F^{↓}R comes with two projections P, Q which are strong arrows (1.3) and a cell ω, called a comma cell

\xymatrix{F^{↓}R \ar[rr]^{P} \ar[dd]|{•}_{Q} &  & A \ar[dd]|{•}^{F} \\  & ω \\ X \ar[rr]_{R} &  & C}  (29)

satisfying the following three universal properties:

(a) (horizontal property) For every pair of horizontal arrows S: U → A, T: U → X and every cell α: (1 (S, RT) F) there is a unique horizontal arrow L such that P L = S, QL = T and moreover the commutative cell λ: (QL = T) gives (λ | ω) = α

\xymatrix{U \ar[rr]^{S} \ar[dd]|{•}_{1^{•}} &  & A \ar[dd]|{•}^{F} &  & U \ar[rr]^{L} \ar[dd]|{•}_{1^{•}} &  & F^{↓}R \ar[rr]^{P} \ar[dd]|{•}^{Q} &  & A \ar[dd]|{•}^{F} \\  & α &  & = &  & λ &  & ω \\ U \ar[r]_{T} & X \ar[r]_{R} & C &  & U \ar[rr]_{T} &  & X \ar[rr]_{R} &  & C}  (30)

(b) (vertical property) For every pair of vertical arrows G: V ^{•} → X, H: V ^{•} → A and every cell β: (G (1, R) F H) there is a unique vertical arrow M such that P M = H, QM = G and the commutative cell μ: (H = P M) gives μ⊗ω = β

[figure: V; •; G; V; •; H; V; •; M; V; •; H; μ; β; A; •; F; =; F^{↓}R; P; •; Q; A; •; F; ω; X; R; C; X; R; C]  (31)

(c) (global property) Suppose we have four arrows L, L′, M, M′, as in the left diagram below. Form the commutative cells λ, μ, and let α = (λ | ω), β = μ⊗ω (as in the previous properties)

[figure: W; L′; •; M′; V; •; M; V; •; H; •; •; •; •; •; •; •; •; •; •; •; σ; μ; α′; U; L; •; 1^{•}; F^{↓}R; P; •; Q; A; •; F; =; •; •; 1^{•}; •; •; •; =; •; •; 1^{•}; β′; •; •; β; •; •; λ; ω; α; U; T; X; R; C; •; •; •; •; •; •]  (32)

then, for every pair of cells (α′, β′) as above, linked by the coherence equation α′⊗α = (β′ | β) (displayed at the right), there is precisely one cell σ: (M′ (L′, L) M) such that:

\begin{matrix}σ | μ = α′,  σ⊗λ = β′ \\ (α′⊗α = (β′ | β), α′ ∈ [M′ (L′, P L) H], β′ ∈ [M′ (L′, T) QM]). \end{matrix}  (33)

We say that 𝔻 has comma objects when all comma objects F^{↓}R exist, for every vertical F and horizontal R having the same codomain.

To study the relationship of comma objects with double limits, it will be useful to consider two particular instances of the global property:

(d) we shall speak of the horizontal 2-dimensional property when we restrict M to be the vertical identity of F^{↓}R; and of the vertical 2-dimensional property when we restrict L to be the horizontal identity of F^{↓}R.

It is easy to see, in the examples below, that the uniqueness of λ (or μ) fails if we do not require that it be a commutative cell: one can start with a non-commutative cell λ′ with the same boundary and replace it with the commutative λ such that (λ | ω) = (λ′ | ω).

3.3. Comments. . A comma object is a sort of ‘symmetric double limit’, related with the horizontal double limits dealt with in Part I and with the symmetrical notion of a vertical double limit, by transpose duality (the symmetry with respect to the main diagonal of double cells).

In fact, by flipping the strong arrow Q: F^{↓}R ^{•} → C to its horizontal companion, the cell ω (29) ‘becomes’ a cell ζ = (η_{Q} | ω)

[figure: F^{↓}R; P; •; 1^{•}; A; •; F; F^{↓}R; •; 1^{•}; F^{↓}R; P; •; Q; A; •; F; ζ; =; η_{Q}; ω; F^{↓}R; RQ; C; F^{↓}R; Q; X; R; C]  (34)

which is (or trivially amounts to) a horizontal cone for the diagram (F, R) in 𝔻, as defined in Part I.

Symmetrically, by flipping the strong arrow P: F^{↓}R ^{•} → A to its vertical companion, the cell ω becomes a cell ϑ

[figure: F^{↓}R; •; 1^{•}; F^{↓}R; •; P; F^{↓}R; •; 1^{•}; F^{↓}R; •; F P; η_{P}; ϑ; =; F^{↓}R; P; •; Q; A; •; F; X; R; C; ω; X; R; C]  (35)

which is a vertical cone of the diagram (F, R) in 𝔻 (as defined by transpose duality).

We shall prove, in Thm. 3.6, that the horizontal (resp. vertical) universal property of the comma is equivalent to the 1-dimensional universal property of the horizontal (resp. vertical) double limit of the diagram (F, R). Furthermore, the global universal property 3.2(c) is a symmetric way of ‘upgrading’ both properties to dimension 2. Thus, a comma object as defined above is at the same time a horizontal and a vertical double limit (in the full, 2-dimensional sense), linked by companionship of projections. Of course, requiring that both such limits exist and coincide is a very strong condition, and we can only expect our comma objects to exist in very particular double categories, having some sort of symmetry in themselves, like 𝔻bl and the double categories of quintets (cf. 3.4). We also note that, without assuming 𝔻 to be strict, the vertical properties should be stated in a much more complex way.

We end with remarking that it would be possible to deal with right Kan extensions using just vertical comma objects, defined as vertical double limits or, equivalently, by the vertical property 3.2(b) together with the vertical 2-dimensional property in 3.2(d) (cf. Thm. 3.6). Symmetrically, left Kan extensions would be dealt with horizontal comma objects, defined as horizontal double limits. However, since we need both aspects and since both kind of limits coincide wherever we need them, it seems simpler to assume this coincidence from the beginning.

3.4. Examples. (a) In the double category of quintets ℚCat, we get the usual comma category F^{↓}R, with objects (a, x; c: F a → Rx), where a is an object of A, x an object of X and c a map of C.

The first two universal properties work in the usual way, with functors L, M defined as follows, on the objects:

L(u) = (Su, T u; αu: F Su → RT u)  (for u in U),
M(v) = (Hv, Gv; βv: F Hv → RGv)  (for v in V). (36)

But the global universal property, in this context, is interesting, and perhaps new. The natural transformation σ: ML′ → LM′ must have components σ(w) = (α′(w), β′(w))

\xymatrix{HL′w \ar[d]_{α′w} & GL′w \ar[d]_{β′w} & F HL′w \ar[r]^{βL′w} \ar[d]_{F α′w} & RGL′w \ar[d]^{Rβ′w} \\ SM′w & T M′w & F SM′w \ar[r]_{αM′w} & RT M′w}  (37)

Conversely, defining σ this way is legitimate: the square above commutes precisely because of the coherence condition α′⊗α = (β′ | β). Finally:

(σ | μ)(w) = P σ(w) = α′w,  (σ⊗λ)(w) = Qσ(w) = β′w.  (38)

(b) Comma objects in 𝔻bl have been constructed in II.2.5 and will be a crucial tool of Part IV. Here, we only recall that a colax double functor F and a lax double functor R with the same codomain have a comma pseudo double category F^{↓↓}R, forming a cell ω in 𝔻bl

[figure: F^{↓↓}R; P; •; Q; 𝔸; •; F; F A; c; •; F a; RX; •; Rx; F A; c; •; F u; RX; •; Rv; ω; =; γ; 𝕏; R; ℂ; F A′; c′; RX′; F B; d; RY]  (39)

An object of the pseudo double category F^{↓↓}R is a triple (A, X; c: F A → RX); a horizontal map (a, x): (A, X; c) → (A′, X′; c′) comes from a commutative square of ℂ, as in the middle diagram above; a vertical arrow (u, v; γ): (A, X; c) ^{•} → (B, Y; d) comes from a cell γ: (F u (c, d) Rv) in ℂ, as in the right diagram above. The projections P, Q are strict double functors (and strong arrows in 𝔻bl).

The first two universal properties of 3.2 have been proved in Thm. II.2.6. The global property can be easily verified, as for ℚCat above, noting that the coherence of the double cell σ can be easily deduced from the coherence of its ‘projections’ on 𝔸 and 𝕏, which are the given cells (σ | μ) = α′, σ⊗λ = β′.

In all these examples, one can note that - coherently with the relationship already mentioned - the comma object is constructed as a (horizontal) double limit, with elements already known from the previous Parts: first the tabulator of the vertical arrow, then a pullback with the given horizontal arrow; finally, one has to take the vertical companion of the ‘second’ projection of this pullback.

3.5. Lemma. [Commutative cells are well determined] Suppose we have, in the weak double category 𝔻, a commutative diagram GR = SF of four horizontal arrows. If all these arrows are strong, the commutative cell λ on the following boundary is well determined: the two constructions considered in 3.1, by flipping the horizontal or the vertical arrows, give the same result

[figure: •; R; •; F_{∗}; •; •; G_{∗}; λ; •; S; •]  (40)

provided we use, in identifying G_{∗}R_{∗} = S_{∗}F_{∗}, the special isocell given by the equality GR = SF, via the units and counits of companionship (see the diagram below).

Proof. The proof is trivial, once the statement is correctly understood. The special isocell referred to above is constructed as follows:

[figure: •; •; 1^{•}; •; •; R_{∗}; •; •; R_{∗}; η_{R}; 1; •; R; •; 1^{•}; •; •; 1^{•}; •; •; G_{∗}; 1^{•}; η_{G}; •; R; F; •; F_{∗}; \begin{matrix}• \\ • \end{matrix}; G; S; •; 1^{•}; •; •; 1^{•}; ϵ_{F}; 1^{•}; •; •; S_{∗}; •; S; •; S_{∗}; •; •; 1^{•}; 1; ϵ_{S}; •; •; •]  (41)

Note that this quaternary vertical composition is uniquely determined, because of a remark in 1.1 on ‘normal composition’: in each of its ‘ternary parts’ there is at least one vertical arrow which is an identity.

3.6. Theorem. [Comma objects and double limits] A comma object F^{↓}R is at the same time the horizontal and the vertical double limit of the diagram (F, R), where the projections are companions.

More precisely, let us consider a cell ω in the double category 𝔻, where P, Q are strong arrows

[figure: Z; P; •; Q; A; •; F; Z; P; •; 1^{•}; A; •; F; Z; •; Q; Z; •; F P; ω; ζ; ϑ; X; R; C; Z; RQ; C; X; R; C]  (42)

and let us write ζ = (η_{Q} | ω) and ϑ = η_{P} ⊗ω the ‘equivalent’ cells obtained by flipping Q or P, as in (34)-(35).

(a) The object Z, equipped with the cell ω, satisfies the horizontal property of comma objects if and only if it satisfies the 1-dimensional universal property of the horizontal double limit of (R, F), when equipped with the horizontal arrows P, Q and the cell ζ. (a′) The object Z, equipped with the cell ω, satisfies the vertical property of comma objects if and only if it satisfies the 1-dimensional universal property of the vertical double limit of (R, F), when equipped with the vertical arrows P, Q and the cell ϑ.

(b) Assume that ζ satisfies the 1-dimensional universal property, or equivalently that ω satisfies the horizontal property. Then, adding the 2-dimensional universal property of ζ is equivalent to adding the horizontal 2-dimensional property of ω (cf. 3.2(d)). (b′) Assume that ϑ satisfies the 1-dimensional universal property, or equivalently that ω satisfies the vertical property. Then, adding the 2-dimensional universal property of ϑ is equivalent to adding the vertical 2-dimensional property of ω (cf. 3.2(d)).

Proof. It is sufficient to prove (a) and (b).

(a) Indeed, the horizontal universal property (30) can be rewritten without commutative cells, saying that

\xymatrix{U \ar[rr]^{S} \ar[dd]|{•}_{1^{•}} &  & A \ar[dd]|{•}^{F} &  & U \ar[rr]^{L} \ar[dd]|{•}_{1^{•}} &  & F^{↓}R \ar[rr]^{P} \ar[dd]|{•}^{1^{•}} &  & A \ar[dd]|{•}^{F} \\  & α &  & = &  & 1^{•} &  & ζ \\ U \ar[r]_{T} & X \ar[r]_{R} & C &  & U \ar[rr]_{L} &  & F^{↓}R \ar[rr]_{RQ} &  & C}  (43)

- for every pair of horizontal arrows S: U → A, T: U → X and every cell α: (1 (S, RT) F) there is a unique horizontal arrow L such that P L = S, QL = T and (1L^{•} | ζ) = α. Which is equivalent to saying that F^{↓}R is the one-dimensional horizontal double limit of the diagram (F, R) in 𝔻, as defined in Part I.

(b) Let us begin by writing down the 2-dimensional property of ζ as a horizontal limit, in a simplified form which takes advantage of the 1-dimensional property, already considered.

A cell σ as below

\xymatrix{U′ \ar[rr]^{L′} \ar[dd]|{•}_{H} &  & F^{↓}R \ar[rr]^{P} \ar[dd]|{•}^{1^{•}} &  & A \ar[dd]|{•}^{F} \\  & σ &  & ζ \\ U′′ \ar[rr]_{L′′} &  & F^{↓}R \ar[rr]_{RQ} &  & C}  (44)

produces two cells

π = (σ | 1P^{•}): (H (P L′, P L′′) 1A^{•}),  ρ = (σ | 1Q^{•}): (H (QL′, QL′′) 1X^{•}),  (45)

satisfying the following relations with the cells ζ′ = (1L′^{•} | ζ), ζ′′ = (1L′′^{•} | ζ) which determine the horizontal arrows L′, L′′:

π⊗ζ′′ = ζ′⊗ (ρ | 1R^{•}),  (46)

since they both coincide with (σ | ζ).

The 2-dimensional universal property says that this mapping is bijective. Precisely, given three arrows and two cells satisfying the following bonds

\begin{matrix}L′: U′ → F^{↓}R, & L′′: U′′ → F^{↓}R, & H: U′ ^{•} → U′′, \\ π: (H (P L′, P L′′) 1A^{•}), & ρ: (H (QL′, QL′′) 1X^{•}), & π⊗ζ′′ = ζ′⊗(ρ | 1R^{•}), \end{matrix}  (47)
\xymatrix{U′ \ar[rrrr]^{P L′} \ar[dd]|{•}_{H} &  &  &  & A \ar[dd]|{•}^{1^{•}} &  & U′ \ar[rr]^{L′} \ar[dd]|{•}_{1^{•}} &  & F^{↓}R \ar[rrr]^{P} \ar[dd]|{•}_{1^{•}} &  &  & A \ar[dd]|{•}^{F} \\  &  & π &  &  &  &  & 1^{•} &  & ζ \\ U′′ \ar[rr]^{L′′} \ar[dd]|{•}_{1^{•}} &  & F^{↓}R \ar[rr]^{P} \ar[dd]|{•}^{1^{•}} &  & A \ar[dd]|{•}_{F} & = & U′ \ar[rr]^{L′} \ar[dd]|{•}_{H} &  & F^{↓}R \ar[r]^{Q} & X \ar[rr]^{R} \ar[dd]|{•}_{1^{•}} &  & C \ar[dd]|{•}^{1^{•}} \\  & 1^{•} &  & ζ &  &  &  &  & ρ &  & 1^{•} \\ U′′ \ar[rr]_{L′′} &  & F^{↓}R \ar[rr]_{RQ} &  & C &  & U′′ \ar[rrr]_{QL′′} &  &  & X \ar[rr]_{R} &  & C}

there is precisely one cell σ: (H (L′, L′′) 1^{•}) satisfying (45), i.e.:

(σ | 1P^{•}) = π,  (σ | 1Q^{•}) = ρ.  (48)

It will be useful to remark that the coherence condition expressed in the diagram above can be equivalently written as follows, by flipping the (strong) projection Q in the right-hand part (ϵ_{Q} is its counit)

π⊗ζ′′ = (ρ′ | ω),  (ρ = (ρ′ | ϵ_{Q})),  (49)
\xymatrix{U′ \ar[rr]^{P L′} \ar[dd]|{•}_{H} &  & A \ar[dd]|{•}^{1^{•}} \\  & π &  &  & U′ \ar[rr]^{L′} \ar[dd]|{•}_{H} &  & F^{↓}R \ar[rr]^{P} \ar[dd]|{•}_{Q} &  & A \ar[dd]|{•}^{F} \\ U′′ \ar[rr]^{P L′′} \ar[dd]|{•}_{1^{•}} &  & A \ar[dd]|{•}^{F} & = &  & ρ′ &  & ω \\  & ζ′′ &  &  & U′′ \ar[rr]_{QL′′} &  & X \ar[rr]_{R} &  & C \\ U′′ \ar[rr]_{RQL′} &  & C}

Now, always by flipping the projections P, Q, we can translate the data of (47) into the following ones, arranged as in the global property of ω, and actually as in the horizontal 2-dimensional property, because of the presence of the vertical identity of F^{↓}R :

[figure: U′; L′; •; H; F^{↓}R; •; 1^{•}; F^{↓}R; •; P; •; L′; •; H; •; •; •; P; •; L′; •; H; •; •; 1^{•}; •; •; P; ?; μ; π′; U′′; L′′; •; 1^{•}; F^{↓}R; P; •; Q; A; •; F; =; •; •; 1^{•}; •; •; •; F; =; •; •; 1^{•}; ρ′; •; •; Q; ϑ; •; •; F; λ; ω; ζ′′; U′′; QL′′; X; R; C; •; QL′′; •; R; •; •; QL′′; •; R; •]  (50)

where

- λ: (QL′′ = QL′′) and μ: (P = P) are commutative cells,

- ζ′′ = (λ | ω) = (1L′′^{•} | ζ), as above, and ϑ = μ⊗ω,

- π′: (H (L′, P L′′) P) corresponds to π: (H (P L′, P L′′) 1A^{•}), by flipping P,

- ρ′: (H (L′, QL′′) Q) corresponds to ρ: (H (QL′, QL′′) 1X^{•}), by flipping Q (as in (48)).

The equivalence of the two universal properties follows now from the following two remarks:

- by flipping P, the coherence condition π⊗ζ′′ = (ρ′ | ω) of (49) is equivalent to π′⊗ζ′′ = (ρ′ | ϑ), i.e. the coherence condition of the global property (50),

- similarly, the conditions on σ in (48), i.e. (σ | 1P^{•}) = π and (σ | 1Q^{•}) = ρ, are equivalent to the conditions for the global property (50), i.e. (σ | μ) = π′ and σ⊗λ = ρ′.

4. Pointwise Kan extensions

Pointwise extensions in a strict double category 𝔻 are defined, using the comma objects of the previous section.

4.1. The pointwise case. (a) Assume that 𝔻 has all comma objects (3.2). We say that G is the pointwise right Kan extension of F, from R to S, via ϵ, if

\xymatrix{H^{↓}R \ar[rr]^{P} \ar[dd]|{•}_{Q} &  & J′ \ar[dd]|{•}^{H} \\  & ω \\ I \ar[rr]^{R} \ar[dd]|{•}_{F} &  & J \ar[dd]|{•}^{G} \\  & ϵ \\ X \ar[rr]_{S} &  & A}  (51)

- for every vertical arrow H: J′ ^{•} → J, GH = Ran_{P,S}(F Q), via ω⊗ϵ, (where ω is the comma cell of H^{↓}R). Then, G is indeed a Kan extension, as we prove below.

(b) We also want a slightly more general definition, where 𝔻 need not have all comma objects. Let 𝔻_{0} be a double subcategory of 𝔻 containing all its objects, all its horizontal arrows, some distinguished vertical arrows (closed under identities and composition, of course) and all the cells whose vertical arrows are distinguished.

Then, if 𝔻_{0} has comma objects, we say that G = Ran_{R,S}(F) is pointwise on the distinguished vertical arrows of 𝔻 if the previous condition holds for every distinguished vertical arrow H: J′ ^{•}→ J. Note that F and G need not be distinguished: the present generalisation does not amount to pointwise Kan extensions in 𝔻_{0}. The fact that G is indeed a Kan extension still holds.

4.2. Theorem. In a pseudo double category 𝔻 with comma objects, every pointwise Kan extension is a Kan extension. More generally, this fact also holds for a Kan extension which is pointwise on a choice of distinguished vertical arrows of 𝔻, in the sense of 4.1(b).

Proof. The proof is based on the Replacement Lemma (2.7). Take, in definition 4.1, J′ = J and H = 1J^{•} (which is necessarily distinguished, according to the previous definition). By hypothesis, G = Ran_{P,S}(F Q). Now, the vertical universal property of J^{↓}R gives a vertical arrow M and a (commuative) cell μ such that

\xymatrix{I \ar[rr]^{R} \ar[dd]|{•}_{M} &  & J \ar[dd]|{•}^{1^{•}} \\  & μ &  & P M = R, \\ J^{↓}R \ar[rr]^{P} \ar[dd]|{•}_{L} &  & J \ar[dd]|{•}^{1^{•}} & LM = 1I^{•}, \\  & ω &  & μ⊗ω = 1R^{•}. \\ I \ar[rr]_{R} &  & J}  (52)

Applying 2.7(a), the fact that G = Ran_{P,S}(F Q) via ω⊗ϵ implies that G = Ran_{R,S}(F) via ϵ.

4.3. Theorem. [Adjunctions are pointwise extensions] Let 𝔻 be a pseudo double category with comma objects. Then an absolute right Kan extension G = Ran_{R,1}(1_{A}), as in Theorem 2.5, is always a pointwise extension.

Proof. With the notation of 2.5, consider the solid left diagram below (omitting G′ and τ), where G is vertically adjoint to R, with cells ϵ, η

[figure: H^{↓}R; P; •; Q; J; •; H; J; •; G′; J; •; GH; J; •; 1^{•}; J; •; 1^{•}; ω; μ; A; R; •; 1^{•}; X; •; G; τ; =; H^{↓}R; P; •; Q; J; •; H; =; 1_{H} ⊗η.; ϵ; ω; A; A; A; A; A; R; X]  (53)

In this situation, the cell η: (G (1, R) 1) yields, with the vertical universal property of the comma, a unique commutative cell μ such that μ⊗ω = 1_{H} ⊗η (see the right diagram above).

A second cell σ: (1^{•} (P, 1) L) comes from the global universal property of ω (3.2(c)), with ω′ = ω⊗ϵ and η′ = 1_{H} ⊗η (since the right-hand equality below plainly holds)

[figure: H^{↓}R; P; •; 1^{•}; J; •; L; J; •; 1^{•}; •; P; •; 1^{•}; •; •; •; 1^{•}; •; P; •; 1^{•}; •; •; •; •; 1^{•}; σ; μ; 1^{•}; H^{↓}R; •; Q; H^{↓}R; P; •; Q; J; •; H; =; •; •; Q; •; •; •; H; =; •; •; Q; ω′; •; •; η′; •; •; H; 1_{Q}; ω; ω; A; A; R; X; •; •; R; •; •; •; R; •]  (54)

(One can note that we are only using the vertical 2-dimensional property, cf. 3.2(d)). We have thus:

(σ | μ) = 1P^{•} ,  σ⊗1_{Q} = ω⊗ϵ.  (55)

Now, given G′: J ^{•} → A and α: (Q (P, 1) G′) on the outer perimeter of the left diagram (53), the existence of a cell τ: (GH (1, 1) G′) such that (ω⊗ϵ | τ) = α determines it, as τ = μ⊗α

μ⊗α = μ⊗(ω⊗ϵ | τ) = (μ⊗ω⊗ϵ | τ) = (1_{H} ⊗η⊗ϵ | τ) = τ,  (56)

where both ternary vertical compositions, μ⊗ω⊗ϵ and 1_{H} ⊗η⊗ϵ, are normal (1.1). Conversely, letting τ = μ⊗α, the problem is solved using (55):

(ω⊗ϵ | τ) = (ω⊗ϵ | μ⊗α) = (σ⊗1_{Q} | μ⊗α) = (σ | μ)⊗α = α.  (57)

5. Pasting properties of comma objects

Pasting properties are proved in Thm. 5.3; as a consequence, pointwise extensions are ‘stable under composition’ (Thm. 5.4).

5.1. Vertical retracts. Recalling some classical facts will give a better understanding of comma objects and their properties. In homotopy theory, a comma square corresponds to a standard homotopy pullback (determined by a similar universal property, where cells are homotopies); diagrammatic lemmas for the latter generally amount to saying that some ‘comparison map’ is a homotopy equivalence. For instance, this is the case for a pasting of homotopy pullbacks; or between the ordinary pullback of a fibration and its homotopy pullback. In the 2-category Cat, one obtains future retracts (i.e., full reflective subcategories) or past retracts (i.e., full coreflective subcategories), according to the direction of cells (cf. [8], 1.6).

Here, we shall obtain somewhat similar results. We say that the horizontal arrow T is a vertical deformation retract of the horizontal arrow R, via the cells λ, μ, φ (as in the diagram below) if we have

\begin{matrix}λ⊗μ = 1T^{•} & & (ML = 1, M′L′ = 1), \\ (φ | μ⊗λ) = 1R^{•}, & 1_{L}⊗φ = 1_{L} & (L′M′ = 1). \end{matrix}  (58)

[figure: •; T; •; L; •; •; L′; •; •; 1^{•}; •; R; •; M; •; •; M′; λ; μ; •; R; •; M; •; •; M′; = 1T^{•} ,; φ; •; T; •; L; •; •; L′; = 1R^{•},; μ; λ; •; T; •; •; •; R; •]

In particular, we shall often encounter the case where R, T are strong arrows, L′ = M′ = 1 and λ, μ are commutative cells. For this case, it is sufficient to have:

R = T M, ML = 1,  1_{L}⊗φ = 1_{L}, φ⊗1_{M} = 1_{M}.  (59)

Then T = T ML = RL, and the commutative cells λ: (T = RL), μ: (R = T M) give

λ⊗μ = 1T^{•} and (φ | μ⊗λ) = φ⊗1_{R} = φ⊗1_{M} ⊗1_{T} = 1_{R}.

5.2. Lemma. [Replacement Lemma, II] In a strict double category 𝔻, two cells ϵ, λ are given, as in the left diagram below (and as in the Replacement Lemma 2.7)

\xymatrix{I′ \ar[rr]^{T} \ar[dd]|{•}_{L} &  & J′ \ar[dd]|{•}^{L′} \\  & λ \\ I \ar[rr]^{R} \ar[dd]|{•}_{F} &  & J \ar[dd]|{•}^{G} \\  & ϵ \\ X \ar[rr]_{S} &  & A}  (60)

If the horizontal arrow T is a vertical deformation retract of the horizontal arrow R, via the cells λ, μ, φ (as in (58)), then the following two conditions are equivalent:

G = Ran_{R,S}(F) via ϵ,  GL′ = Ran_{T,S}(F L) via λ⊗ϵ.  (61)

Proof. It is a straightforward consequence of the previous Replacement Lemma, 2.7.

5.3. Theorem. [Pasting Theorem] Let 𝔻 be a strict double category. Consider the pasting of two comma cells, in the left diagram, and the comma ω′′ of the vertical composite F F′, in the middle diagram

[figure: F′^{↓}P; P′; •; Q′; B; •; F′; (F F′)^{↓}R; P′′; •; Q′′; B; •; F F′; •; •; 1^{•}; F′^{↓}P; P′; •; M; B; •; 1^{•}; ω′; μ; F^{↓}R; P; •; Q; A; •; F; ω′′; φ; (F F′)^{↓}R; P′′; •; L; B; •; 1^{•}; ω; λ; X; R; C; X; R; C; •; F′^{↓}P; P′; B]  (62)

Then the strong arrow P′′ (a projection of the comma of the composite) is a vertical deformation retract (5.1) of P′ (a projection of the iterated comma), in the strong sense of (59):

- there exist commutative cells λ, μ and a comparison cell φ, as displayed in the right diagram above, satisfying:

ML = 1,  1_{L}⊗φ = 1_{L},  φ⊗1_{M} = 1_{M}.  (63)

Note. Recall that ML = 1 gives λ⊗μ = 1 (and actually amounts to that, since our cells are commutative).

Proof. First, by the vertical universal property of ω′′, there is a unique vertical arrow M: F′^{↓}P ^{•} → (F F′)^{↓}R such that

P′′M = P′,  Q′′M = QQ′,  μ⊗ω′′ = ω′⊗ω,  (64)
\xymatrix{F′^{↓}P \ar[rr]^{P′} \ar[dd]|{•}_{M} &  & B \ar[dd]|{•}^{1^{•}} &  & F′^{↓}P \ar[rr]^{P′} \ar[dd]|{•}_{Q′} &  & B \ar[dd]|{•}^{F′} \\  & μ &  &  &  & ω′ \\ (F F′)^{↓}R \ar[rr]^{P′′} \ar[dd]|{•}_{Q′′} &  & B \ar[dd]|{•}^{F F′} & = & F^{↓}R \ar[rr]^{P} \ar[dd]|{•}_{Q} &  & A \ar[dd]|{•}^{F} \\  & ω′′ &  &  &  & ω \\ X \ar[rr]_{R} &  & C &  & X \ar[rr]_{R} &  & C}

with the commutative cell μ: (P′ = P′′M).

Now, to define the second comparison L, we begin by constructing a vertical arrow N: (F F′)^{↓}R ^{•} → F^{↓}R, by the vertical universal property of F^{↓}R (3.2(b))

P N = F′P′′,  QN = Q′′,  ν⊗ω = ω′′ (ν: (F′P′′ = P N)),  (65)
[figure: (F F′)^{↓}R; •; N; (F F′)^{↓}R; •; F′P′′; ν; (F F′)^{↓}R; P′′; •; N; B; •; F′; F^{↓}R; P; •; Q; A; •; F; = ω′′,; ν; ω; F^{↓}R; P; A; X; R; C]

Flipping horizontally the strong arrow P′′, the cell ν will also be written as at the right hand, above.

Then, we have a vertical L: (F F′)^{↓}R ^{•}→ F′^{↓}P, by the vertical universal property of F′^{↓}P

P′L = P′′,  Q′L = N,  λ⊗ω′ = ν: (F′P′′ = P N).  (66)
\xymatrix{(F F′)^{↓}R \ar[rr] \ar[dd]|{•}_{L} &  & B \ar[dd]|{•}^{1^{•}} \\  & λ \\ F′^{↓}P \ar[rr] \ar[dd]|{•}_{Q′} &  & B \ar[dd]|{•}^{F′} & = ν \\  & ω′ \\ F^{↓}R \ar[rr]_{P} &  & A}

The equality λ⊗μ = 1P′′^{•} (i.e., ML = 1) is then detected by the vertical universal property of (F F′)^{↓}R

(λ⊗μ)⊗ω′′ = λ⊗ω′⊗ω = ν⊗ω = ω′′.  (67)

As a first step for constructing φ, the global universal property of ω gives a cell ρ satisfying the following equations, where \overline{ω} = ω′ ⊗ω (notice that P′ and Q should be ‘flipped’, to follow the pattern of 3.2(c))

[figure: •; •; Q′; F′^{↓}P; P′; •; M; B; •; 1^{•}; μ; ρ; (F F′)^{↓}R; P′′; •; N; B; •; F′; •; •; Q′; •; P′; •; •; F′; •; •; Q′; •; P′; •; •; •; F′; ν; ω′; •; •; Q; F^{↓}R; P; •; Q; A; •; F; =; •; •; Q; •; •; •; F; =; •; •; Q; 1; •; •; \overline{ω}; •; •; F; 1; ω; ω; •; X; R; C; •; •; R; •; •; •; R; •]  (68)

(ρ | (μ⊗ν)) = ω′,  ρ⊗1_{Q} = 1_{QQ′}.  (69)

It will be useful to note that 1_{L}⊗ρ = 1_{N}, as it results applying λ⊗− = (1_{L} | λ)⊗− to all terms of diagram (68). In fact, λ⊗μ = 1, λ⊗ω′ = ν and λ⊗(ω′⊗ω) = ν⊗ω = ω′′.

[figure: •; •; N; (F F′)^{↓}R; P′′; •; N; B; •; F′; •; •; N; •; P′′; •; •; F′; •; •; Q′; •; P′′; •; •; •; F′; 1_{L}⊗ρ; ν; ν; •; •; Q; F′^{↓}P; P; •; Q; B; •; F; =; •; •; Q; •; •; •; F; =; •; •; Q; 1; •; •; ω′′; •; •; F; 1; ω; ω; •; X; R; C; •; •; R; •; •; •; R; •] (70)

Since putting 1_{N} in the upper-left cell above also solves these equations, these solutions coincide.

The global universal property of ω′ gives a cell φ satisfying the following conditions, with μ⊗ν = ζ (we know from (68) that (ρ | ζ) = ω′)

[figure: •; •; 1^{•}; F′^{↓}P; P′; •; M; B; •; 1^{•}; μ; φ; (F F′)^{↓}R; P′′; •; L; B; •; 1^{•}; •; •; 1^{•}; •; P′; •; •; 1^{•}; •; •; 1^{•}; •; P′; •; •; •; 1^{•}; λ; 1^{•}; •; •; Q′; F′^{↓}P; •; Q′; B; •; F′; =; •; •; Q′; •; •; •; F′; =; •; •; Q′; ρ; •; •; ζ; •; •; F′; 1; ω′; ω′; •; F^{↓}R; P; A; •; •; P; •; •; •; P; •] (71)

φ | (μ⊗λ) = 1P′^{•},  φ⊗1_{Q′} = ρ.  (72)

Now, the equalities 1_{L} ⊗φ = 1_{L} and φ⊗1_{M} = 1_{M} of the thesis are detected by the global universal properties of ω′ and ω′′. The first comes from the fact that, applying λ⊗− = (1_{L} | λ)⊗− on the top of diagram (71), we get

λ⊗μ⊗λ = λ,  1_{L}⊗ρ = 1_{N},  λ⊗μ⊗ν = ν,  (73)

[figure: •; •; L; (F F′)^{↓}R; P′′; •; L; B; •; 1^{•}; •; •; L; •; P′′; •; •; 1^{•}; •; •; L; •; P′′; •; •; •; 1^{•}; 1_{L}⊗φ; λ; λ; •; •; Q′; F′^{↓}P; •; Q′; B; •; F′; =; •; •; Q′; •; •; •; F′; =; •; •; Q; 1; •; •; ω′′; •; •; F; 1; ω′; ω′; •; F^{↓}R; P; A; •; •; P; •; •; •; P; •] (74)

and these equations are also solved by putting 1_{L} in the upper-left cell above.

The second comes from a similar uniqueness argument applied to the following equations:

[figure: •; •; M; (F F′)^{↓}R; P′′; •; M; B; •; 1^{•}; •; •; M; •; P′′; •; •; 1^{•}; •; •; M; •; P′′; •; •; •; 1^{•}; φ⊗1_{M}; μ; μ; •; •; Q′′; (F F′)^{↓}R; •; P′′; B; •; F F′; =; •; •; Q′′; •; •; •; =; •; •; Q′′; 1; •; •; \overline{ω}; •; •; 1; ω′′; ω′′; •; X; R; C; •; •; R; •; •; •; R; •] (75)

which are also solved by 1_{M}.

5.4. Theorem. [Pointwise stability] Let 𝔻 be a strict double category. If G = Ran_{R,S}(F) is a pointwise right Kan extension, then for every vertical arrow H: J′ ^{•} → J the extension GH = Ran_{P,S}(F Q) (see the diagram below) is still a pointwise extension, via the cell ω⊗ϵ.

Proof. Follows from the first Replacement Lemma (2.7) together with the Pasting Theorem 5.3.

Take an arbitrary H′: K′ ^{•} → K and its comma H′^{↓}P, as in the left diagram below

\xymatrix{H′^{↓}P \ar[rr]^{P′} \ar[dd]|{•}_{Q′} &  & J′′ \ar[dd]|{•}^{H′} & H′^{↓}P \ar[rr]^{P′} \ar[dd]|{•}_{L} &  & J′′ \ar[dd]|{•}^{1^{•}} \\  & ω′ &  &  & λ \\ H^{↓}R \ar[rr]^{P} \ar[dd]|{•}_{Q} &  & J′ \ar[dd]|{•}^{H} & (HH′)^{↓}R \ar[rr]^{P′′} \ar[dd]|{•}_{Q′′} &  & J′′ \ar[dd]|{•}^{HH′} \\  & ω &  &  & ω′′ \\ I \ar[rr]_{R} \ar[dd]|{•}^{F} &  & J \ar[dd]|{•}^{G} & I \ar[rr]_{R} &  & J \\  & ϵ \\ X \ar[rr]_{S} &  & A}  (76)

We have to prove that (GH)H′ = Ran_{P′,S}(F QQ′), via ω′⊗(ω⊗ϵ). Forming the comma (HH′)^{↓}R, in the right diagram, the pointwise property of G with respect to the composite HH′ says that

G(HH′) = Ran_{P′′,S}(F Q′′), via ω′′⊗ϵ.  (77)

Now, the Pasting Theorem 5.3 says that the projection P′′ is a vertical deformation retract of P′, with a commutative comparison cell λ such that λ⊗ω′′ = ω′⊗ω. Applying the Replacement Lemma 2.7(b) to (2), we get the thesis: (GH)H′ = Ran_{P′,S}(F QQ′), via λ⊗(ω′′⊗ϵ) = ω′⊗ω⊗ϵ.

6. Kan extensions for double categories

We briefly examine now the case when 𝔻 is a double category of double categories. One of such settings will be studied in Part IV, together with its relations with double limits.

6.1. External Kan extensions. In a (pseudo) double category 𝔻, we have defined a right Kan extension G = Ran_{R,S}(F) via ϵ, as in the left diagram below (see 2.2)

[figure: 𝕀; R; •; F; 𝕁; •; G; 𝕀; R; •; F; 𝕁; •; G; ϵ; η; 𝕏; S; 𝔸; 𝕏; S; 𝔸]  (78)

and a left Kan extension S = Lan_{F,G}(R) via η, as in the right diagram above (2.3).

These notions will be used letting 𝔻 be one of the three ‘settings’ for weak double categories listed below. Thus, the vertices 𝕀, 𝕁, 𝕏, 𝔸 will be pseudo double categories, all arrows will be some kind of ‘weak’ double functors, and ϵ or η some kind of horizontal transformation GR → SF.

6.2. Colax right Kan extensions. For 𝔻 = 𝔻bl (1.4), the general definition G = Ran_{R,S}(F) says that G is the colax right Kan extension of the colax double functor F, along the lax double functors R, S (or from R to S), via the cell ϵ (the counit of the extension). The pointwise case is defined as in 4.1, since 𝔻bl has all comma objects (3.4(b)).

Symmetrically, S = Lan_{F,G}(R) means that S is the lax left Kan extension of a lax double functor R, along the colax double functors F, G, via the cell η (the unit of the extension).

We have already seen that, in a colax/lax adjunction G ⊣ R between weak double categories, the left adjoint is an absolute right Kan extension of this type, G = Ran_{R,1}(1) (2.5), and is always a pointwise extension (4.3). Symmetrically, the right adjoint is an absolute left Kan extension R = Ran_{1,G}(1) (since transpose duality exchanges left and right adjoints).

Note that, here, transposing 𝔻 is equivalent to applying ‘internal’ horizontal duality to the vertices, arrows and cells of diagram (78); in fact the latter reverses the direction of cells and exchanges lax with colax double functors.

6.3. Unitary colax right Kan extensions. Let 𝔻bl_{u} be the cell-wise full double subcategory of 𝔻bl where the vertical arrows are unitary colax double functors, while the horizontal ones are general. By restriction, also 𝔻bl_{u} has all comma objects (since in 𝔻bl the projections of a comma are strict double functors).

Taking 𝔻 = 𝔻bl_{u} gives a unitary colax right Kan extension of a unitary colax double functor F: 𝕀 ^{•}→ 𝕏, along two lax double functors. The interest of restricting vertical arrows to be unitary is shown by a few examples below (7.2, 7.3).

The symmetric notion we are interested in, unitary lax left Kan extensions along colax double functors, lives in 𝔻 = 𝔻bl^{u}, where the horizontal arrows are restricted to be unitary (lax). Again, internal horizontal duality on all data exchanges these situations.

6.4. Unitary lax right Kan extensions. . Finally, we write Lx𝔻bl_{u} the double category of weak double categories, lax double functors (horizontally) and unitary lax double functors (vertically); a cell α: GR → SF, as in the diagram above, simply is a horizontal transformation of the composed lax double functors.

(Thus, Lx𝔻bl_{u} is a substructure of the double category of quintets Lx𝔻bl = ℚLxDbl, over the 2-category LxDbl of weak double categories, lax double functors and their horizontal transformations, already considered in II.2.2; the whole Lx𝔻bl is not of interest for the present extensions, see 7.4.) Taking 𝔻 = Lx𝔻bl_{u} gives a unitary lax right Kan extension of a unitary lax double functor F: 𝕀 ^{•} → 𝕏, along two lax double functors. But here, we can always compose SF obtaining an (arbitrary) lax double functor, and reduce our data to the (equivalent) case where the left arrow is an identity

[figure: 𝕀; R; •; 1^{•}; 𝕁; •; G; ϵ; 𝕀; S; 𝔸]  (79)

Now, Lx𝔻bl_{u} does not have all comma objects, but its double subcategory Lx𝔻bl_{p} = 𝔻bl_{p} (with unitary pseudo double functors as vertical arrows) does: they are again a part of comma objects in 𝔻bl. Speaking of pointwise Kan extensions in the present setting we will always mean pointwise on unitary pseudo double functors, as defined in 4.1(b).

This framework is perhaps the most adequate for studying pointwise right extensions and their relationship with double limits, and will be dealt with in Part IV.

Here, the symmetric notion of interest, unitary colax left Kan extensions along colax double functors, lives in 𝔻 = Cx𝔻bl^{u}, where the horizontal arrows are unitary colax double functors and the vertical ones are colax. Again, it comes from internal horizontal duality on all data. This setting is adequate for studying pointwise left extensions and their relationship with double colimits,

7. Examples

We end with some examples showing how the different settings we have considered in the last section can influence the resulting Kan extension.

7.1. Remarks. Some important cases arise when 𝕁 is the terminal double category \underline{1}, or the formal vertical arrow \underline{2}, i.e. the double category which has one vertical arrow 0 ^{•} → 1 and is otherwise trivial.

Note that a unitary lax (or colax) double functor defined on \underline{1} or \underline{2} is necessarily strict, as already remarked at the end of 1.1. Therefore, unitary lax or colax double functors defined on 𝕁 = \underline{1} or \underline{2} coincide, and the ‘colax case’ G = Ran_{R,S}(F) (displayed below, at the left) is more general than the ‘lax’ one, G = Ran_{R,S}(1) (displayed at the right, as in 6.4)

[figure: 𝕀; R; •; F; 𝕁; •; G; 𝕀; R; •; 1^{•}; 𝕁; •; G; ϵ; ϵ; 𝕏; S; 𝔸; 𝕀; S; 𝔸]  (80)

7.2. Double limits as pointwise unitary (co)lax extensions. By the previous remarks, the following situation can equivalently be viewed in 𝔻bl_{u} or Lx𝔻bl_{u}.

Let A = lim(S) be the (horizontal) double limit of a lax double functor S: 𝕀 → 𝔸, as defined in Part I.

It is easy to see that this amounts to saying that the unitary double functor A: \underline{1} → 𝔸 is the pointwise unitary (colax or lax) right Kan extension of S along the projection R: 𝕀 → \underline{1}, via the limit cone ϵ: AR → S (see the right diagram above, with 𝕁 = \underline{1}).

On the other hand, a lax (resp. colax) double functor T: \underline{1} → 𝔸 is a monad (resp. a comonad) in the bicategory V𝔸 of objects, vertical arrows and special cells of 𝔸, and the universal such can be quite different from the limit, as shown below (7.3, 7.4).

One can note that the unitary colax extension G = Ran_{R,S}(F), as in the left diagram (80), would give the ‘generalised double limit’ of a composite SF: 𝕀 ^{•} → 𝕏 → 𝔸, where the colax F need not be the identity. These generalised double limits can still be constructed from the elementary limits of Part I; but the relevance of this fact is not clear, and we shall not develop it here.

7.3. Unitary colax versus colax. Comparing the setting 𝔻bl_{u} with 𝔻bl, it is easy to construct examples where the unitary colax Kan extension on the singleton (the limit) is different from the ‘’general’ colax Kan extension, and actually more interesting.

Let 𝔸 = ℚCat, the double category of quintets on the 2-category Cat, so that V𝔸 is the cell-dual of Cat, with reversed cells.

An ordinary monad (A, T, η, μ) can be viewed as a strict double functor T: \underline{m} → 𝔸, where \underline{m} is the formal vertical monoid, i.e. the strict double category generated by a vertical arrow t: 0 ^{•} → 0, with two special cells e: 1 → t, m: t^{2} → t linked by the monoid axioms.

Its non-unitary colax Ran on the projection \underline{m} → \underline{1} coincides with T, viewed as a colax double functor T: \underline{1} → 𝔸 (a comonad in V𝔸 and a monad in Cat). But its double limit, i.e. the unitary (colax or lax) Ran (see 7.1), is the category of algebras A^{T} , with cone based on the forgetful functor

\begin{matrix}ϵ_{0} = U^{T} : A^{T} → A, & ϵ_{1_{0}} = ϵ: T U^{T} → U^{T} : A^{T} → A, \\ ϵ(A, a: T A → A) = a: T A → A. \end{matrix}  (81)

Indeed, let us apply our construction theorem for double limits (I.5.5) to T: \underline{m} → 𝔸. One begins (I.6.6) by replacing the vertical arrows U = 1A^{•}, T, T^{2} with their tabulators and the natural transformations η, μ (viewed as cells of ℚCat) with the corresponding arrows

[figure: 𝔸; •; T; 𝔸; •; U; 𝔸; •; T; 𝔸; •; T^{2}; η; μ; ⊤U; ⊤T; ⊤η; ⊤μ; ⊤T^{2}; 𝔸; 𝔸; 𝔸; 𝔸]  (82)

where ⊤T has objects (A, B, f: T A → B); similarly for ⊤U and ⊤T^{2}; and

\begin{matrix}⊤η: ⊤T → ⊤U, & (⊤η)(A, B, f: T A → B) = (A, B, f.ηA: A → B), \\ ⊤μ: ⊤T → ⊤T^{2}, & (⊤μ)(A, B, f: T A → B) = (A, B, f.μA: T^{2}A → B). \end{matrix}  (83)

Then, as explained in I.6.7, one adds the following arrows d, p, q, c, to take into account the fact that U is a vertical identity and T^{2} a vertical composite (the objects of the ‘iterated tabulator’ ⊤(T, T) are specified below)

\xymatrix{⊤U & ⊤T \ar[l]_{⊤η} \ar[r]^{⊤μ} & ⊤T^{2} \\ 𝔸 \ar[u]^{d} &  & ⊤(T, T) \ar[u]_{c} \ar@{-->}[ul]^{p} \ar@{-->}[ul]_{q}}  (84)

d(A) = (A, A, 1_{A}),

p(A, B, C; f: T A → B; g: T B → C) = (A, B, f), q(A, B, C; f: T A → B; g: T B → C) = (B, C, g), c(A, B, C; f: T A → B; g: T B → C) = (A, C, g.T f: T^{2}A → C).

Finally, the double limit of T is the ordinary limit of the diagram (84). It is a full subcategory of ⊤T (because d is monic, and the pair p, q jointly monic); and it contains precisely those objects (A, B, f: T A → B) of ⊤T which are algebras

B = A,  f.ηA = 1_{A};  f.μA = f.T f.  (85)

7.4. Unitary lax versus lax. One can give a similar, simpler example to compare the setting Lx𝔻bl_{u} with Lx𝔻bl.

With 𝕀 = \underline{1}, again, a lax double functor S: \underline{1} → ℚCat^{h} (with values in the horizontally opposite double category) amounts to a monad on the category S(0). Its double limit A, corresponding to the unitary right Kan extension, is the category of Eilenberg-Moore algebras, while the general lax right Kan extension is trivial and coincides with S.

7.5. Local products. The following works both in 𝔻bl_{u} and Lx𝔻bl_{u}. Take for 𝕀 the strict double category \underline{2}_{I}, having two objects, 0 and 1, and a small set I of vertical arrows i: 0 ^{•} → 1. Let R: \underline{2}_{I} → \underline{2} be the obvious projection, with R(i) = (0 ^{•} → 1)

[figure: \underline{2}_{I}; R; •; 1^{•}; \underline{2}; •; G; ϵ; _{\underline{2}_{I}}; u; 𝔸]  (86)

A unitary (necessarily strict) double functor u: \underline{2}_{I} → 𝔸 is a family of parallel vertical arrows u_{i}: A ^{•}→ B. The pointwise unitary, lax or colax, Kan extension G = Ran_{R,u}(1) amounts to an arrow u: A_{0} ^{•}→ A_{1} equipped with a family of special cells σ_{i}: u → u_{i}, universal in the obvious sense. We shall call it the local product of the family.

This does not exist in ℚSet or ℚCat (showing that this extension is independent of the existence of double limits). But it does exist in ℝel (the double category of sets, mappings and relations, see I.3.4), where it gives the intersection of our relations (even for a large I). It also exists (for a small I) in 𝕊et, where it is constructed with a limit. And in ℂat, where we have

\begin{matrix}u: A^{op}×B → Set, & u(a, b) = ∏_{i} u_{i}(a, b), \\ p_{i}: u → u_{i}, & p_{i}(a, b): u(a, b) → u_{i}(a, b). \end{matrix}  (87)

7.6. Lax versus colax. This example is an extension of the previous one. It is based on the strict double category \underline{3}, which is again a vertical ordinal, generated by two vertical arrows 0 ^{•}→ 1 ^{•}→ 2. Of course, unitary lax or colax double functors \underline{3} → 𝔸 no longer coincide.

Take as 𝕀 the (vertical) double category freely generated by three objects 0, 1, 2 and two sets of vertical arrows

u′_{i}: 0 ^{•} → 1,  u′′_{j}: 1 ^{•} → 2  (i ∈ I, j ∈ J).  (88)

It has thus vertical arrows u′′_{j}u′_{i}: 0 ^{•} → 2 indexed on I ×J, and is horizontally discrete. Form now the following diagram

[figure: 𝕀; R; •; 1^{•}; \underline{3}; •; G; ϵ; 𝕀; S; 𝔸]  (89)

where R is the obvious projection (preserving objects), 𝔸 = ℝel, and S is a strict double functor.

First, let us consider the setting 𝔻 = Lx𝔻bl_{u} and let G be the unitary lax right Kan extension. Plainly, G is computed by intersection of parallel relations (the ‘local products’ of 7.5), and is a pointwise extension

\begin{matrix}G(0 ^{•} → 1) = ⋂_{i} F (u′_{i}), & G(1 ^{•} → 2) = ⋂_{j} F (u′′_{i}), \\ G(0 ^{•} → 2) = ⋂_{ij} F (u′′_{i}u′_{i}), & G(0 ^{•} → 1)⊗G(1 ^{•} → 2) ⊂ G(0 ^{•} → 2). \end{matrix}  (90)

Note that G is not strict (i.e., not colax), generally: one can easily construct a finite example where both G(0 ^{•} → 1) and G(1 ^{•} → 2) are empty, while G(0 ^{•} → 2) is not.

On the other hand, the unitary colax right Kan extension, in (89), also exists, and has G′(0 ^{•} → 2) = G(0 ^{•} → 1)⊗G(1 ^{•} → 2); therefore, it is not pointwise.

References

[1] J. Bénabou, Introduction to bicategories, in: Reports of the Midwest Category Seminar, Lecture Notes in Mathematics, Vol. 47, Springer, Berlin 1967, pp. 1-77.

[2] R. Brown and G. Mosa, Double categories, 2-categories, thin structures and connections, Theory Appl. Categ. 5 (1999), 163-175.

[3] R. Bruni, J. Meseguer and U. Montanari, Symmetric monoidal and Cartesian double categories as a semantic framework for tile logic, Math. Structures Comput. Sci. 12 (2002), 53-90.

[4] R.J.M. Dawson, R. Paré and D.A. Pronk, Paths in double categories, Theory Appl. Categ. 16 (2006), 460-521.

[5] C. Ehresmann, Catégories structurées, Ann. Sci. Ecole Norm. Sup. 80 (1963), 349-425.

[6] C. Ehresmann, Catégories et structures, Dunod, Paris 1965.

[7] T.M. Fiore, Pseudo algebras and pseudo double categories, J. Homotopy Relat. Struct. 2 (2007), no. 2, 119-170.

[8] M. Grandis, The shape of a category up to directed homotopy, Theory Appl. Categ. 15 (2005/06), No. 4, 95-146.

[9] M. Grandis and R. Paré, Limits in double categories, Cah. Topol. Géom. Différ. Catég. 40 (1999), 162-220.

[10] M. Grandis and R. Paré, Adjoint for double categories, Cah. Topol. Géom. Différ. Catég. 45 (2004), 193-240.

[11] P.A. Melliès, Double categories: a modular model of multiplicative linear logic, Math. Structures Comput. Sci. 12 (2002), no. 4, 449-479.

[12] S.S. Moskaliuk and T.A. Vlassov, Double categories in mathematical physics, International Symposium on Mathematical and Theoretical Physics (Kyiv, 1997), Ukraïn. Fīz. Zh. 43 (1998), no. 6-7, 836-841.

[13] C.B. Spencer and Y.L. Wong, Pullback and pushout squares in a special double category with connection, Cah. Topol. Géom. Différ. 24 (1983), 161-192.

[14] M. Shulman, Framed Bicategories and Monoidal Fibrations, Preprint 2007, arXiv:0706.1286.

[15] R. Street, Fibrations and Yoneda’s lemma in a 2-category, in: Category Seminar, Sydney 1972/73, Lecture Notes in Mathematics, Vol. 420, Springer, Berlin 1974, pp. 104-133.

Dipartimento di Matematica Università di Genova Via Dodecaneso 35 16146-Genova, Italy

Department of Mathematics and Statistics Dalhousie University Halifax NS Canada B3H 3J5

Email: grandis@dima.unige.it
      pare@mathstat.dal.ca

This article may be accessed at http://www.tac.mta.ca/tac/ or by anonymous ftp at

ftp://ftp.tac.mta.ca/pub/tac/html/volumes/20/8/20-08.{dvi,ps,pdf}

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods. Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication. Full text of the journal is freely available in .dvi, Postscript and PDF from the journal’s server at http://www.tac.mta.ca/tac/ and by ftp. It is archived electronically and in printed paper format.

Subscription information. Individual subscribers receive abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh, rrosebrugh@mta.ca.

Information for authors. The typesetting language of the journal is TEX, and L^{A}TEX2e strongly encouraged. Articles should be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files at http://www.tac.mta.ca/tac/.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: barr@math.mcgill.ca

Assistant TEX editor. Gavin Seal, Georgia Southern University: gseal@georgiasouthern.edu

Transmitting editors.

Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown (at) btinternet.com Aurelio Carboni, Università dell Insubria: aurelio.carboni@uninsubria.it Valeria de Paiva, Xerox Palo Alto Research Center: paiva@parc.xerox.com Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk P. T. Johnstone, University of Cambridge: ptj@dpmms.cam.ac.uk Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, University of Western Sydney: s.lack@uws.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@acsu.buffalo.edu Jean-Louis Loday, Université de Strasbourg: loday@math.u-strasbg.fr Ieke Moerdijk, University of Utrecht: moerdijk@math.uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Brooke Shipley, University of Illinois at Chicago: bshipley@math.uic.edu James Stasheff, University of North Carolina: jds@math.unc.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Insubria: robert.walters@uninsubria.it R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca